\newpage
\begin{thebibliography}{10}

\bibitem{bala_ranga}
R.~Balakrishnan and K.~Ranganathan, \emph{A {T}extbook of {G}raph {T}heory},
  Springer, New York, 2000.

%\bibitem{bbm94}
%{Y.~A.} Ban, S.~Bereg, and {N.~H.} Mustafa, \emph{A {C}onjecture on {W}iener
%  {I}ndices in {C}ombinatorial {C}hemistry}, Algorithmica \textbf{40} (2004),
%  99--117.

%\bibitem{CK97}
%V.~Chepoi and Klav\v{z}ar, \emph{The {W}iener {I}ndex and the {S}zeged {I}ndex
%  of {B}enzenoid {S}ystems in {L}inear {T}ime}, J.~Chem.~Inf.~Comput.~Sci.
%  \textbf{37} (1997), 752--755.

%\bibitem{dank93}
%P.~Dankelmann, \emph{Computing the {A}verage {D}istance of an {I}nterval
%  {G}raph}, Inf.~Proc.~Lett. \textbf{48} (1993), 311--314.

\bibitem{dobry1}
{A.~A.} Dobrynin, R.~Entringer, and I.~Gutman, \emph{Wiener {I}ndex of {T}rees:
  {T}heory and {A}pplications}, Acta {A}ppl. {M}ath. \textbf{66} (2001), 211--249.
  
\bibitem{dobry2}
{A.~A.} Dobrynin, I.~Gutman, S.~Klav\v{z}ar and P.~\v{Z}igert, \emph{Wiener {I}ndex of {H}exagonal
  {S}ystems}, Acta {A}ppl. {M}ath. \textbf{72} (2002), 247--294.

\bibitem{epp99}
D.~Eppstein, \emph{Subgraph {I}somorphism in {P}lanar {G}raphs and {R}elated
  {P}roblems}, J.~Graph Algo.~Appl. \textbf{3(3)} (1999), 1--27.

\bibitem{es2001}
E.~Estrada and E.~Uriarte, \emph{Recent {A}dvances on the {R}ole of
  {T}opological {I}ndices in {D}rug {D}iscovery {R}esearch}, Cur.~Med.~Chem.
  \textbf{8} (2001), 1573--1588.

\bibitem{polansky}
I.~Gutman and O.~Polansky, \emph{Mathematical {C}oncepts in {O}rganic
  {C}hemistry}, Springer-{V}erlag, Berlin, 1986.

\bibitem{ms05}
M.~Miller and J.~Siran, \emph{Moore {G}raphs and {B}eyond: {A} {S}urvey of the
  {D}egree / {D}iameter {P}roblem}, The Elec.~J.~Combin. \textbf{DS14} (2005),
  1--61.

%\bibitem{wag06}
%{S.~G.} Wagner, \emph{A {C}lass of {T}rees and its {W}iener {I}ndex}, Acta
%  Appl. Math. \textbf{91(2)} (2006), 119--132.

\bibitem{wiener}
H.~Wiener, \emph{Structural {D}etermination of {P}araffin {B}oiling {P}oints},
  J. Amer. Chem. Soc. \textbf{69} (1947), 17--20.

\bibitem{xu}
L.~Xu and X.~Guo, \emph{Catacondensed {H}exagonal {S}ystems with {L}arge
  {W}iener {N}umbers}, MATCH {C}ommun. {M}ath. {C}omput. {C}hem. \textbf{55}
  (2006), 137--158.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage{enumerate,amsmath,amsthm,mathrsfs,array,pstricks, amssymb}
\allowdisplaybreaks[1]

\newtheoremstyle{bthm}{\baselineskip}{\baselineskip}{\slshape}{}{\bfseries}{:}{\newline }{}
\newtheoremstyle{bex}{\baselineskip}{\baselineskip}{}{}{\sffamily}{:}{\newline }{}
\theoremstyle{bthm}
\newtheorem{thm}{Theorem}

\newtheorem{rem}{Remark}
\begin{document}
\title{A sharp lower bound for the Wiener index of  a graph}
\author{R. Balakrishnan\\ Srinivasa Ramanujan Centre\\Kumbakonam--612 001, India\\ 
email: mathbala@sify.com 
\and
N. Sridharan\\
Department of Mathematics\\
Alagappa University\\
Karaikudi--630 003, India\\
email: math\_sridhar@yahoo.co.in
\and  K. Viswanathan Iyer\\
Department of Computer Science and Engg.\\ National Institute of Technology\\ 
Trichy--620 015, India\\
email: kvi@nitt.edu
}
\date{}
\maketitle
\renewcommand{\baselinestretch}{1.58}\small\normalsize
\begin{abstract}
Given a simple connected undirected graph $G$, the Wiener index $W(G)$ of $G$ is defined as half the sum of the distances over all pairs of vertices of $G$.  In practice, $G$ corresponds to what is known as the \emph{molecular graph} of an  organic compound.   We obtain a sharp lower bound for $W(G)$ of an arbitrary graph  in terms of the order, size and  diameter of $G$.  
\end{abstract}
\section{Introduction}
Let $G=\big(V(G),\,E(G)\big)$ be a simple  connected undirected graph of order $n$ and size $m$. Given any two vertices $u,\,v$ of $G$, let  $d(u,\, v)$ denote the distance between $u$ and $v$. 
The Wiener index $W(G)$ of the graph $G$ is defined by
\[ W(G)=\frac{1}{2}\sum_{u, v\in V(G)}d(u,\, v),\]
where the summation is over all possible pairs $u,\, v \in V(G)$.  Our notation and terminology are as in~\cite{bala_ranga}.

Wiener index, first proposed in \cite{wiener}, is currently a widely used topological index (see, for example, \cite{xu}) and has applications in modern drug design~\cite{es2001}.  For more information about Wiener index in chemistry and in mathematics, see~\cite{polansky}, \cite{dobry1} and \cite{dobry2}. In this paper, for any graph $G$, we obtain a sharp lower bound for $W(G)$ in terms of the order, size and diameter of $G$.

\section{Lower bound for $W(G)$}
Let $S_2(G)$ be the set of all 2-subsets of $V(G)$ (that is, the set of all unordered pairs of distinct vertices of $G$).  We can then equivalently define $W(G)$ as
\begin{equation}
W(G)=\sum_{\{u,v\}\in S_2(G)}d(u,v).\label{bound}
\end{equation}
Let $P=u_0u_1\ldots u_d$ be a diametral path of $G$, so that $d(u_0,u_d)=d$, the diameter of $G$.  We partition $S_2(G)$ into disjoint sets $X,Y$ and $Z$ defined as follows:
\begin{align*}
X&=\{\{u,v\}\in S_2(G)\vert \quad \text{both}\  u,v\in P\},\\
Y&=\{\{u,v\}\in S_2(G)\vert \quad \text{none of $u$ and $v$ is in $P$}\},\quad\text{and}\\
Z&=\{\{u,v\}\in S_2(G)\vert \quad \text{one of $u$ and $v$ alone is in $P$}\}.
\end{align*}
It follows that
\[
\vert X\vert =\frac{d(d+1)}{2};\quad \vert Y\vert =\frac{(n-d-1)(n-d-2)}{2};\quad \vert Z\vert =(n-d-1)(d+1).
\]
From~\eqref{bound}, we have
\begin{align}
W(G)&=\sum_{\{u,v\}\in S_2(G)}\big(2+(d(u,v)-2)\big)\nonumber\\
&=\sum_{\{u,v\}\in S_2(G)}2\; + \sum_{\{u,v\}\in S_2(G)}\big(d(u,v)-2\big)\label{wpapeqn12}\\
&= n(n-1) 
+\sum\limits_{\substack{\{u,\,v\}\in S_2 (G)\\d(u,\,v)=1}} \big(d(u,\,v)-2\big) + \mathop{\sum}\limits_{\substack{\{u,\,v\}\in S_2 (G)\\d(u,\,v)\geq 2}} \big(d(u,\,v)-2\big)\nonumber\\
\intertext{(since $\vert S_2(G)\vert  = n(n-1)/ 2)$}
&= \big(n(n-1) -m\big) + \mathop{\sum}\limits_{\substack{\{u,\,v\}\in S_2(G)\\d(u,\,v)\geq 2}} \big(d(u,\,v)-2\big)\nonumber\\
&\ge (n(n-1) - m) + \sum_{\substack{\{u,\,v\} \in X\cup Z\\d(u,\,v)\ge2}} \big(d(u,\,v)-2\big)\label{wpapeqn13}.
\end{align}
For $0\le k \le(d-1)$ in $X$, there are $(d-k)$ pairs $\{u,\,v\}$ with $d(u,\,v) = 1 + k$.  Hence
\begin{align}
\sum_{\substack{\{u,\,v\} \in X\\d(u,\,v)\ge 2}} \big(d(u,\,v) - 2\big) &= (d-2)1 + (d-3)2 + \cdots + 1(d-2)\nonumber\\
							 &= \frac{d(d-1)(d-2)}{6}. 	
\end{align}
We next obtain a lower bound for the summation term on the right hand side of equation~\eqref{wpapeqn13}.  We first assume that $d\ge 5$.  Fix one vertex $w$ in $V(G)\setminus V(P) $, where $V(P)$ is the set of vertices of $P$.  Then, by triangle inequality, we have
\begin{align}
d(u_i,\,w) + d(w,\,u_{d-i})\ge \; &d(u_i,\,u_{d-i}) = d-2i,\\ 
&\text{for}\quad 0\le i < (d-3)/2.\nonumber
\end{align}
Therefore, for each of the $(n-d-1)$ choices of $w$, we have
\begin{align}
\sum_{d(u_i,w) \ge 2} \big(d(u_i,\,w)-2\big) \nonumber \\
\quad \ge \sum_{i=0}^d \big(d(u_i,\,w)-2\big)
&\ge \sum_{i=o}^{\lfloor\frac{d-3}{2}\rfloor} (d(u_i,\,w) + d(u_{d-i},\,w)-4)\label{wpapeqn16}\\
&\ge \sum_{i=o}^{\lfloor\frac{d-3}{2}\rfloor} (d-2i-4).\label{wpapeqn17}
\end{align} 
Since each summand on the right side of~\eqref{wpapeqn16} is nonnegative,  so is each summand on the right side of~\eqref{wpapeqn17}.  Hence the term on the right side of~\eqref{wpapeqn16} is
\begin{align*}
& \ge 
\begin{cases}
(d-4)+(d-6) +\cdots + 5+3+1 \quad \text{if $d$ is odd, and}\\
(d-4) + (d-6) + \cdots + 2 \quad \text{if $d$ is even}.
\end{cases}\\[2ex]
&= 
\begin{cases}
(\frac{d-3}{2})^2 \quad \text{if $d$ is odd, and}\\
\frac{(d-2)(d-4)}{4}\quad \text{if $d$ is even}.
\end{cases}
\end{align*}
Thus, for each fixed $w \in V(G) \setminus V(P)$ we have 
\[
\sum_{(u,\,w)\in Z} (d(u,\,w)-2) = 
\begin{cases}
(\frac{d-3}{2})^2 \quad \text{if} \; d \;\text{is odd, and}\\
\frac{(d-2)(d-4)}{4}\quad\text{if} \; d\;\text{is even}. 
\end{cases}
\]
In conclusion, for $d \ge 5$, we have
\[
W(G) \geq 
\begin{cases}
(n(n-1)-m + \frac{d(d-1)(d-2)}{6} + \frac{(n-d-1)(d-3)^2}{4},\quad \text{if $n$ is odd}, \\
(n(n-1)-m) + \frac{d(d-1)(d-2)}{2} + \frac{(n-d-1)(d-2)(d-4)}{4}, \; \text{if $n$ is even}. 
\end{cases}
\]
We now consider the cases when $d = 2,\,3\,$ and 4.
If $d=2$, $d(u,\,v) = $ 1 or 2. Hence from~\eqref{wpapeqn12}, we get 
\begin{align}
W(G) &= n(n-1) + \sum_{d(u,\,v)=1} (1-2)\nonumber\\
		 &= n(n-1)-m.
\end{align}
Now consider the case when $d = 3$. Then $d(u,\,v) = $1,\,2 or 3.  Hence from~\eqref{wpapeqn12} we get
\begin{align*}
W(G) &= n(n-1)-m + \sum_{d(u,\,v)=3} (d(u,\,v)-2)\\
&\ge n(n-1)-m + 1,
\end{align*}
as there is at least one pair with $d(u,\,v)=3$.  If $d = 3$, $\dfrac{d(d-1)(d-2)}{6} = 1$.  Hence we have 
\[
W(G) \ge n(n-1)-m + \frac{d(d-1)(d-2)}{6}.
\]
Finally if $d = 4$, we have 
\begin{align*}
W(G) \ge n(n-1)-m  &+ \sum_{d(u,\,v) =3} (d(u,\,v)-2) \\ &+ \sum_{d(u,\,v)=4} (d(u,\,v)-2)\\
&\geq\;n(n-1)-m+2+2\tag*{(since $d(u_0,\,u_d)=d= 4$, $d(u_0,\,u_3)=2=d(u_1,\,u_4)$)}.
\end{align*}
Hence
\begin{align*}
W(G)&\geq n(n-1)-m+\frac{d(d-1)(d-2)}{6} \tag*{\big(\text{as $\frac{d(d-1)(d-2)}{6}=4$ in this case.\big)}}.
\end{align*}
To conclude, we have proved the following result:
\begin{thm}
\label{sharp-thm-1}
If $G$ is any graph of order $n$, size $m$  and diameter $d\ge 2$ then
\begin{align}
W(G)\geq
\begin{cases}
n(n-1)-m+\frac{d(d-1)(d-2)}{6}+\frac{(n-d-1)(d-3)^2}{4},&\text{if $n$ is odd.}\\
n(n-1)-m+\frac{d(d-1)(d-2)}{6}+\frac{(n-d-1)(d-2)(d-4)}{4},&\text{if $n$ is even.}\label{new-9}
\end{cases}
\end{align}
\end{thm}
\begin{rem}
Let the maximum degree of $G$ be $\Delta$.  The Moore bound (see for example~\cite{ms05}) gives an upper bound for $n$ in terms of $\Delta$ and the diameter $d$:
\begin{equation}
n\leq 
\begin{cases}
1+ \Delta \frac{(\Delta -1)^d-1}{\Delta -2}& \text{if $\Delta >2.$}\\
2d+1 & \text{if $\Delta=2.$}
\end{cases}\label{new-10}
\end{equation}
\noindent
From~\eqref{new-10}, a lower bound for $d$ in terms of $n$ and $\Delta$ can be obtained and if this is used in~\eqref{new-9}, a lower bound for $W(G)$ can be obtained in terms of $n$, $m$ and $\Delta$.
\end{rem}
\begin{rem}
If $G$ is known to be planar and if $d$ is known to be bounded by a constant then we can compute the exact value of $d$ in time $O(n)$ as shown in \cite{epp99}.  In turn, this implies that the lower bound on $W(G)$ in~\eqref{new-9} can be computed in time $O(n)$ if $G$ is planar.
\end{rem} 
We observe that the lower bound given in Theorem~\ref{sharp-thm-1} is sharp.  The graphs $P_n$, $K_{1,m}$ ($m$:odd), $C_3\, \Box\, K_2$ and the Petersen graph attain the bound.

\end{document}